OpenStudy (anonymous):

is the point (1, 4, 2) or the point (2, 2, -6) on the graph of 3x -2y +z + 4 =0

OpenStudy (aakashsudhakar):

Equation: 3X - 2Y + Z + 4 = 0 Points to Check: (1, 4, 2) & (2, 2, -6) Simply plug in the values to see if they work! Plug in (1, 4, 2) to get 3(1) - 2(4) + (2) + 4 = 0 And simplify. 3 - 8 + 2 + 4 = 0 -5 + 2 + 4 = 0 -3 + 4 = 0 -1 = 0 ^^ That does not work, so the point (1, 4, 2) does not lie on the original function. Now plug in (2, 2, -6) to get 3(2) - 2(2) + (-6) + 4 = 0 And simplify. 6 - 4 - 6 + 4 = 0 You may already see how this will all cancel out. 2 - 6 + 4 = 0 -4 + 4 = 0 0 = 0 ^^ That does work, so the point (2, 2, -6) lies on the original function.